### v1

# LlamaTouch: A Faithful and Scalable Testbed for Mobile UI Automation Task Evaluation

Abstract.
The emergent large language/multimodal models facilitate the evolution of mobile agents, especially in the task of mobile UI automation. However, existing evaluation approaches, which rely on human validation or established datasets to compare agent-predicted actions with predefined ones, are unscalable and unfaithful. To overcome these limitations, this paper presents LlamaTouch, a testbed for on-device agent execution and faithful, scalable agent evaluation. By observing that the task execution process only transfers UI states, LlamaTouch employs a novel evaluation approach that only assesses whether an agent traverses all manually annotated, essential application/system states. LlamaTouch comprises three key techniques: (1) On-device task execution that enables mobile agents to interact with real mobile environments for task completion. (2) Fine-grained UI component annotation that merges pixel-level screenshots and textual screen hierarchies to explicitly identify and precisely annotate essential UI components with a rich set of designed annotation primitives. (3) A multi-level state matching algorithm that utilizes exact and fuzzy matching to accurately detect critical information in each screen with unpredictable UI layout/content dynamics. LlamaTouch currently incorporates four mobile agents and 495 UI automation tasks, encompassing both tasks in the widely-used datasets and our self-constructed ones for more diverse mobile applications. Evaluation results demonstrate the LlamaTouch’s high faithfulness of evaluation in real environments and its better scalability than human validation. LlamaTouch also enables easy task annotation and integration of new mobile agents. Code and dataset are publicly available at https://github.com/LlamaTouch/LlamaTouch.
copyright: none
1. Introduction
Mobile intelligent agents empower users to interact with their smartphones using natural languages, alleviating them from tedious, cumbersome smartphone operations. These agents are particularly beneficial for individuals with visual or hand impairments, or in situations where using a screen is not practical (e.g., driving). Notable mobile agents, such as Apple Siri (Apple, 2024) and Google Assistant (Google, 2024b), have become indispensable services on smartphones. The recent advent of large language models (LLMs) and multimodal LLMs have facilitated researchers building more powerful mobile agents (Li et al., 2024; Wen et al., 2023; Wang et al., 2023; Lee et al., 2023). The key capability of those agents is to comprehend user instructions in natural language and execute corresponding actions on the mobile interface, as called mobile UI automation, e.g., “forward the last email from Bob to Alice”.
Despite claims of powerful UI automation capabilities achieved by recent LLM-powered mobile agents, their evaluation methods are somewhat flawed. Unlike traditional machine learning models evaluated on well-established, static datasets, mobile agents need to interact with the dynamic, indeterministic states of a smartphone (e.g., network connectivity, dynamic content) as inputs. Simply evaluating mobile agents using deterministic smartphone states from datasets cannot uncover their true capabilities (Rawles et al., 2023; Li et al., 2020; Sun et al., 2022).
In general, there are two methods to evaluate mobile UI automation tasks, but neither achieves both high faithfulness and scalability. (1) The most intuitive approach is to request humans to verify the completion of tasks. However, human evaluation is difficult to reproduce (Chiang and Lee, 2023). The requisite human effort increases with the number of agents, tasks, and evaluation platforms. (2) The most popular approach used in most prior arts (Rawles et al., 2023; Yan et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023; Hong et al., 2023) is exact action match on established datasets, akin to traditional machine learning evaluations. The key idea is to ask annotators to generate a correct sequence of actions that succeed on the task as data labels, and then compare agent-generated actions to these labels. Although this approach allows for some error tolerance, e.g., the variations in click positions on the screen (Rawles et al., 2023), it cannot cover all possible and “infinite” paths to complete a UI automation task. Consequently, it leads to a significantly higher false negative rate. For example, for task “Reserve a rental car in Los Angeles from June 1st-7th, with a budget of up to \$60 per day on Expedia”, the sequence of three filtering actions can be interchanged. Only taking one of these execution paths as the reference may incorrectly verify a task that is essentially completed. Moreover, LLM-powered agents are known to be able to self-correct their wrong actions (Pan et al., 2023), which is critical to enhancing UI automation capabilities, yet is impossible to evaluate in a static dataset. These limitations are further demonstrated in $\S$§$2.$
This paper presents LlamaTouch, the first testbed for evaluating mobile agents on real-world mobile environments without compromising faithfulness and scalability. The key idea of LlamaTouch is to check the task execution trace against a few “essential states” identified by the annotators, rather than matching them against predefined action sequences in static traces. For instance, the essential states for the task “open app Microsoft Excel (install if not already installed), go to login, …” should include (1) the application “Microsoft Excel” is opened; and (2) the application is on the login page. Other operations, like app installation, are considered non-essential and should be ignored. During task execution, LlamaTouch enables mobile agents to retrieve only the task description from the static dataset, while device states are directly acquired from real-world devices. Actions produced by mobile agents are executed on those smartphones, and all UI interaction data are recorded as task execution traces. In the evaluation phase, LlamaTouch compares task execution traces with annotated essential states to determine whether a task has been completed.
To ensure faithful and scalable evaluation, LlamaTouch integrates two effective methods. (1) LlamaTouch adopts a fine-grained labeling mechanism for essential state annotation at both the screen level and single UI component level. It combines pixel-level screenshots and textual screen hierarchies to explicitly highlight important UI components. With a rich set of annotation primitives provided by LlamaTouch, it reduces human efforts to heuristically identify and annotate the attributes of essential states for evaluation, e.g., the content of a textbox should be exactly matched. These annotated primitives are subsequently used for faithful evaluation. (2) During the evaluation process, LlamaTouch employs a multi-level state match algorithm that combines fuzzy and exact matches on diverse annotated UI states. It uses (i) approximate screen match that enables LlamaTouch to adapt to dynamic mobile environments and varying screen contents, and (ii) mixed UI state match that detects and matches critical on-screen information.
Dataset and testbed. We present a large-scale dataset with pre-annotated essential states for evaluating mobile UI automation tasks in real-world Android environments. This dataset includes 495 distinct tasks encompassing a wide array of popular Android applications. We complement this dataset with an easy-to-use testbed that enables mobile agents to interact seamlessly with real Android environments. This testbed provides a collection of concise, widely used APIs, ensuring compatibility with most mobile agents. Mobile agents can be easily integrated into LlamaTouch and use our dataset to test their capabilities in mobile UI automation tasks in real-world scenarios.
Evaluation. We implemented LlamaTouch by utilizing Google Android emulator (Google, 2023) as actual smartphones. Currently, LlamaTouch has four built-in agents, AutoUI (Zhan and Zhang, 2023), AppAgent (Yang et al., 2023), AutoDroid (Wen et al., 2023), and CoCo-Agent (Ma et al., 2024), along with 495 diverse tasks. The number of supported agents and tasks will be extended in the future. With human validation as the ground truth for task completion, LlamaTouch achieves nearly 80% accuracy in detecting completed tasks in real-world environments, while exact action match on static datasets fails to do so. We also reveal the limitations of current mobile agents in handling tasks practically in real-world environments.
Contributions are summarized as follows.
  • We observed the weakness of high false negative rates in evaluating mobile UI automation agents using static datasets. To address this, we proposed an evaluation design that only compares essential states rather than concrete action sequences.
  • We devised a method for annotating essential states using a variety of annotation primitives. This approach combines visually intuitive screenshots with semantically precise VHs to enable fine-grained and accurate UI component localization and annotation.
  • We designed a novel task evaluation system that employs both exact and fuzzy matching at various UI state levels. It enables faithful evaluation of mobile agents and scales well to dynamic execution environments.
  • We proposed LlamaTouch, the first testbed to faithfully and salably evaluate mobile UI automation agents in real-world mobile environments. It comprises 495 UI automation tasks with human-annotated essential states. Four agents integrated in LlamaTouch demonstrate its faithfulness and scalability in UI automation task evaluation.
2. Background and Motivation
2.1. Agents for Mobile UI Automation
Mobile agents have simplified the cumbersome and dull operations on smartphones for users. The progression of mobile agents for mobile UI task automation can be categorized into three phases. (1) API-based agents like Google Assistant (Google, 2024b) and Apple Siri (Apple, 2024) interact with applications through pre-defined application programming interfaces. This approach is reliable while limited in structured and predictable tasks. (2) Learning-based agents (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023) utilize deep learning techniques to learn from previous mobile interaction traces, but their capabilities were still confined by their training data. (3) Recently, LLMs and multi-modal LLMs have revolutionized the capabilities of mobile agents (OpenAI, 2023; Hong et al., 2023; Wen et al., 2023; Wang et al., 2023). These models, owing to their vast knowledge base, can understand complex, real-world scenarios. Mobile agents powered by these models are able to interpret natural language instructions accurately and translate them into actionable tasks on smartphones’ screens. This evolution marks a significant leap in the flexibility and adaptability of mobile agents.
Mobile agents typically operate with the following components. Controller is the brain of mobile agents. It interprets task instructions and UI contexts, and generates actions to be executed on the current UI context. The widely-used controllers include DL models tailored for specific applications (Rawles et al., 2023; Sun et al., 2022; Li et al., 2020; Zhan and Zhang, 2023), LLMs (GPT-3.5, GPT-4) (Wen et al., 2023; Wang et al., 2023; Lee et al., 2023), and multi-modal LLMs (GPT-4V) (Yan et al., 2023; Hong et al., 2023; Yang et al., 2023).
Input: UI representation. Existing mobile agents take a task description and a UI representation as the input of their controller. There are two basic types of UI representations: screenshots and VH. A screenshot is a visual capture of the current screen. A text-format VH provides a tree-like structure of the UI elements present on a screen, including their properties such as type, position, and text contents. On top of screenshots and VH, some controllers further extract UI semantics to enhance UI understanding. For example, Yan et al. (Yan et al., 2023) overlies numeric tags on top of each text and icon detected by OCR tools; AXNav (Taeb et al., 2023) also converts screenshots to bounding boxes and labels, enabling them comprehensible by LLMs. Further processing based on VH, like converting it to simple HTML representations, is also widely utilized (Wang et al., 2023; Wen et al., 2023).
Output: actions. The output of controllers are actions to be executed on the current screen, such as click, swipe, and input text. Action parameters can be abstracted at different levels depending on the agent’s design and their input format. (1) Concrete coordinates on the screen (Rawles et al., 2023; Sun et al., 2022; Hong et al., 2023; Zhan and Zhang, 2023). It operates as a direct interaction with the screen like human operations. (2) Icon marker (Yan et al., 2023; Yang et al., 2023). The output target will specify a specific icon or graphical element within the UI representation. (3) HTML index (Wang et al., 2023; Wen et al., 2023; Lee et al., 2023). By ingesting HTML representations, controllers will give a concrete HTML index as the action target, which matches specific elements (icons or text) on the screen.
Table 1. The comparison between mobile agent benchmarks. LlamaTouch is the first testbed designed for mobile agents driven by essential state match. LlamaTouch also supports fine-grained UI-guided essential state annotation with a rich set of primitives covering a wide array of diverse UI states.
Benchmark Platform
Real-world
Tasks
Real-env
Task Exec
Fine-grained
UI Annotation
Essential
State Match
Rico (Deka et al., 2017) Mobile
PixelHelp (Li et al., 2020)
AndroidEnv (Toyama et al., 2021)
META-GUI (Sun et al., 2022)
MoTIF (Burns et al., 2022)
AITW (Rawles et al., 2023)
Mobile-Env (Zhang et al., 2023)
AndroidArena (Xing et al., 2024)
WebArena (Zhou et al., 2023) Web
LlamaTouch Mobile
Figure 1. Two major limitations of evaluating mobile agents on static datasets.
$2.$ Mobile UI Automation Benchmarks
As shown in Table 1, a variety of datasets and environments are proposed to evaluate mobile agents in UI automation tasks, while none of them achieve both faithfulness and scalability.
Some work such as Rico (Deka et al., 2017), PixelHelp (Li et al., 2020), and AITW (Rawles et al., 2023), provides static datasets (task descriptions and UI representations). Mobile agents predict concrete actions on static UI representations, which will be compared with the ground-truth actions in datasets. While this approach is straightforward, it is insufficient to reveal the performance of mobile agents for two reasons. (1) Inaccurate exact action match are pervasive when a functionally correct action is deemed incorrect due to different action parameters. Figure 1(a) shows two cases. First, for click actions, the clickable area defined in the dataset may be narrow. While in reality, the area might encompass the entire UI component (marked with the red bounding box). Second, non-identical text inputs can lead to the same correct result (marked with green bounding boxes) in most search tasks. This issue has also been observed in previous literature but has not been solved (Yan et al., 2023; Rawles et al., 2023). (2) Lack of tolerance for different execution paths. In real-world environments, an UI automation task can usually be completed in various paths, as demonstrated in Figure 1(b1) and Figure 1(b2). However, pre-defined datasets might only provide one deterministic path for reference, leading to inaccurate evaluation.
There are also work that enable agent execution in real-world environments. AndroidEnv (Toyama et al., 2021) and Mobile-Env (Zhang et al., 2023) enable the execution of mobile automation tasks in real environments. However, they do not essentially support essential state match during the end-to-end task execution, therefore compromising evaluation accuracy. AndroidArena (Xing et al., 2024) observed the weakness of step-wise action match on static datasets does not fully tolerant redundant actions in task execution paths. They proposed a subsequence-based action match: a task is treated as completed when it contains the ground-truth action sequence as its subsequence. We take AndroidArena as a baseline in $\S$§4.3 to compare its evaluation accuracy with LlamaTouch. WebArena (Zhou et al., 2023) provides a real playground for Web agents. It uses essential states to evaluate task completion (e.g., the final result should be or should include some key information). LlamaTouch differs from WebArena on the mobile platform in both the essential state annotation and evaluation process: (i) LlamaTouch combines the visual screenshot with textual view hierarchy of the same screen for fine-grained and precise UI component identification; (ii) LlamaTouch uses a richer set of primitives to comprehensively annotate essential UI states and faithfully evaluate them even with high screen content dynamics.
Human validation is usually used to validate whether a UI automation task is completed (Rawles et al., 2023; Taeb et al., 2023). However, the cost of human validation is too high, thereby scaling poorly to multiple tasks, agents, and mobile devices. LlamaTouch ensures high scalability as evaluating on static datasets while preserving faithfulness as human validation.
Figure 2. The workflow of on-device task execution.
3. LlamaTouch Design
Compared to previous evaluation approaches, LlamaTouch exhibits the following benefits.
$\bullet$ Practical on-device task execution ($\S$§3.1). Most of the previous evaluation methods are simulated on mobile UI interaction datasets. LlamaTouch, on the other side, enables mobile agent evaluation on real mobile devices, revealing their real capabilities in the wild.
$\bullet$ Fine-grained essential application state annotation ($\S$§3.2). By observing UI automation task execution only transfers essential UI elements within an application, LlamaTouch enables annotators to explicitly annotate essential application states that should be detected and matched during task execution.
$\bullet$ Faithful and scalable task evaluation ($\S$§3.3). LlamaTouch evaluates mobile agents by comparing task execution traces with annotated application states. By combining exact matching with fuzzy matching algorithms on different application states, LlamaTouch achieves faithful and scalable task evaluation.
3.1. On-device Task Execution
Evaluating mobile agents on pre-defined traces, as discussed in Section 2.2, depicts significant inaccuracy. LlamaTouch empowers on-device UI automation task execution to reveal the real capabilities of mobile agents. To simplify this process, LlamaTouch operates as a bridge between existing mobile agents and real mobile devices (e.g., smartphones, Android emulators, and cloud device farms). As shown in Figure 2, mobile agents execute tasks on LlamaTouch by taking the following processes. 1 Mobile agents request one task instruction from LlamaTouch. The instruction comes from one supported dataset by LlamaTouch, e.g., AITW. 2 With this task instruction, mobile agents will then request UI representation (e.g., screenshot, VH) from LlamaTouch, which will be 3 forwarded to mobile devices. 4 Taking the task instruction and UI representation as inputs, mobile agents will predict an action to be performed on the current UI, and dispatch predicted actions to LlamaTouch. 5 LlamaTouch forwards and executes the agent-predicted action to real mobile devices. The processes from 2 to 5 are repeatedly executed until mobile agents treat the task as completed. During task execution, all UI representations and corresponding actions are captured for further evaluation in $\S$§$3.$ These UI representations (1) pixel-level screenshots; (2) textual screen VHs; (3) activity names of the application at the foreground of each screen; and (4) actions performed on every screen. Essential system states such as the list of installed applications will also be dumped for faithful mobile agent evaluation, as we will illustrate in $\S$§3.2. Appendix A.1 details APIs provided by LlamaTouch for mobile agent integration.
Table 2. Primitives of essential states in LlamaTouch that are used in the annotation and evaluation processes.
Match Type State Type Primitive Keyword Use Case
Fuzzy match UI state Screen info fuzzy<-1> Check if the contents on two screens are approximately identical.
Textbox fuzzy<n>
Check if the content of the target textbox is semantically similar
to the content of the original textbox<n> in the ground-truth UI.
Exact match Activity activity
A coarse-grained approach to determining if two UIs represent
the same functional screen in an application.
UI component
exact<n>,
exclude<n>
Check if the UI component is exactly identical to the UI component<n>,
or does not occur, in the ground-truth UI.
System state (Un)installation
installed<app>,
uninstalled<app>
Check if the target application named "app" has been successfully
installed/uninstalled.
Action Action
click<n>,
type<input_text>
Check if two actions and their parameters are identical.
Figure 3. Annotated essential state for task “empty the shopping cart on bestbuy” in the last UI representation: two textboxes that should be exactly matched. The essential state represents the application state after task execution: “the shopping cart on bestbuy is empty”.
3.2. Essential Application State Annotation
Two major cases demonstrated in $\S$§2.2 make exact matching on pre-defined traces unfaithful. First, it demands that the agent-generated actions and their parameters match exactly with those in the dataset. Second, it takes a fixed UI interaction sequence provided by the dataset as the reference, therefore unable to evaluate alternative task execution paths.
Insight: the task execution process transfers identifiable application states. The internal of the execution process of a UI automation task is changing application states that some of them can be explicitly represented by UI components. Even if the task execution paths differ, there are overlaps in the essential application states. We can utilize these overlapping application states to determine whether a task achieves some milestones or is completed. As the example shown in Figure 1(b), by identifying the intent of the task “open app Microsoft Excel (install if not already installed), go to login, …”, it contains two potential essential states for evaluation: (1) the application “Microsoft Excel” is opened; and (2) the application is located at the login page. Other actions, like detecting whether the application is installed, can be omitted during evaluation.
To do so, essential states should be accurately identified and annotated. However, simply annotating application states at the whole UI representation level (e.g., a screenshot) and comparing screen-level similarity (Feiz et al., 2022), are far coarse-grained and may lead to inaccurate matching. For example, the screen contents of a web-shopping application could be subtly different due to slight movement or dynamically loaded contents across different executions. For accurate and efficient essential state annotation, LlamaTouch breaks the whole pixel-level UI representation (i.e., screenshot) to separate UI components. LlamaTouch simplifies the textual VH of each screen, which precisely expresses the attributes of every UI element, to extract important, visible UI components. These extracted UI components are combined with visually intuitive screenshots to provide precise overlayed bounding boxes and unique identifiers to annotators. Figure 3 shows an example of the VH-enhanced screenshot provided to annotators.
Primitives. LlamaTouch incorporates a list of primitives for essential state annotation. These primitives are used to extract essential information on the screen. Annotators are instructed to clearly identify and use primitives to represent application states that are informative and deterministic to validate the task execution results. Table 2 comprehensively shows these primitives and their use cases. Overall, LlamaTouch incorporates six types of primitives and nine keywords for annotation. These primitives can be divided into three types according to the state type they represent.
$\bullet$ UI state. The purpose of UI state annotation is to extract and compare whether two screens contain identical or similar information. As a screen may contain different types of components, LlamaTouch has four major primitives to cover all these UI components. (1) Screen info and (2) textbox are used to compare whether the whole screen or a dedicated textbox has similar contents. Screen info is primarily used for checking whether two UI representations are within the same screen (i.e., the same page of an application). Textbox primitive is used to annotate textboxes that may contain dynamic contents, such as a search box in a web-shopping application or the URL of a website. (3) The functionality of activity is like screen info: it can be used to approximately detect whether two UI representations are on the same screen of an application. (4) UI component is used to deterministically annotate the state of a UI component, including textbox, button, image, etc. For example, if the content of a textbox should be identical, it should be annotated with the exact keyword. States that require certain UI components should not occur on the screen should be annotated with the exclude keyword.
$\bullet$ System state. LlamaTouch involves system state annotation mainly inspired by the tasks in a previously built dataset: AITW (Rawles et al., 2023). For example, a task “install app YouTube Kids” can be detected directly using shell commands (e.g., pm list packages), without involving redundant UI operations. Currently, LlamaTouch supports programmatically checking application installation. The keywords can be easily extended when other system states are required to be detected.
$\bullet$ Action. Although the principle of annotation is to detect application state transfer during task execution. Sometimes, action on a specific UI representation is necessary to validate agent behaviour, especially when there is not enough state identifier shown in UI. LlamaTouch provides click and type keywords that covers most common actions.
Case study. Essential state annotation can be divided into the following processes. First, for each UI representation, LlamaTouch overlayed all functional UI components with numeric markers to the screenshot as shown in Figure 3. This is done by extracting the precise metadata of each UI component from textual VH. Then, annotators will identify what are essential states that should be checked during evaluation to ensure the task is completed. LlamaTouch simplify this process by only requiring annotators to explicitly identify what UI element should be matched. For example, after emptying the shopping cart on bestbuy, the screen will display a textbox with the content “Your cart is empty”. This textbox should be treated as an essential state, which is highlighted in a bounding box with numeric id 13 in the last screen of Figure 3, as it represents the state after task completion. As we anticipate the content of the textbox should be exactly matched, the annotated keyword is exact<13>. The keyword exact<27> is used to validate whether two screens are both in the shopping cart of bestbuy. All annotations along with the UI representations and task descriptions construct the essential state-powered dataset. These essential states are used by LlamaTouch for faithful task evaluation.
$3.$ Faithful and Scalable Task Evaluation
Annotated essential states ($\S$§3.2) and task execution traces captured on real devices ($\S$§3.1) are used for evaluation on LlamaTouch. LlamaTouch iterates essential states to determine whether a task execution trace goes through all annotated states sequentially. If so, this task is deemed completed. To achieve faithful evaluation, the most significant challenge lies in how to ensure the task execution trace matches the essential states. Task execution traces captured from real-world environments may contain dynamic screen content; it is vital to adapt the annotated essential states on a static dataset to dynamic, real-world contents, while achieving precise match on only critical information. To address this problem, LlamaTouch employs a multi-level state matching algorithm, which combines fuzzy match and exact match on both the entire screen and separated UI components, to ensure faithful evaluation. It first approximately matches two screens according to their UI representation and activity. Then, within two matched screens, it compares the annotated primitives that should be fuzzily or exactly matched.
Approximate screen match is used to ensure two screens are on the same functionality page of an application, even under high screen content dynamics. This is the prior process for in-screen exact/fuzzy content match. To achieve this, LlamaTouch utilizes two annotation primitives proposed in $\S$§3.2: activity and screen info.
1. Application activity match. Activity represents an entry point for users to interact with the application (and, 2024). This is an obvious identifier to indicate whether two screens are in the same application. Screens with distinct functionalities typically have unique activity names, even if they are within the same application. For example, the activity of the main setting page is com.android.settings.Settings, while the value is com.wifiadmin.settings.WifiSettingsActivity in the Wi-Fi setting page. Typically, exact application activity match acts as a foundational filtering process to identify whether two screens are on the same page. However, some specific application design philosophies may make different functional pages of an application contain the same activity name (sin, 2018). Under this circumstance, LlamaTouch involve screen info for accurate screen match.
2. Fuzzy screen info match is utilized when the activity name cannot reveal different functional pages of an application. To better compare two screens, the primary thing is to only extract critical information from the screen. In summary, LlamaTouch simplifies the textual screen VH to a simple HTML representation as in prior work (Wen et al., 2023; Wang et al., 2023), while containing the types of every single UI component. LlamaTouch compares two simplified HTML representations of two screens with a cosine similarity of their sentence embeddings. If the result falls within a threshold, LlamaTouch treats two screens are matched. Two screens are similar when they are above a predefined threshold, e.g., 0.85 in our experiments. The fuzzy screen info match design helps LlamaTouch maintain faithfulness during evaluation when dealing with dynamic screen contents.
Mixed UI state match will be applied to matched screens after the approximate screen match process. A mix of annotated UI states will be checked in this phase, including both fuzzy match and exact match on UI components, actions, and system states.
1. Exact UI component match requires the annotated UI component should be identical in two screens to be compared, including all their attributes such as class, text, and selected in VH. This is especially useful for evaluating the content of a textbox, the status of a button (e.g., checked or not), or a selectable icon on the screen. Given a target UI component to be exactly matched, LlamaTouch will iterate nodes in VH of the matched screen, until one matched UI component is found. Exact UI component match fails if there is no matched node on available screens. In our dataset, exact UI component match contributes to 50.5% (688 out of 1,362) of annotated essential states.
2. Fuzzy textbox match is crucial for comparing content inside a textbox on the screen, especially when the content may be slightly different. Semantically similar searching keywords with the same intent that complies with a specific task description should be matched, as they will lead to the same results. For example, Figure 1(b) shows searching “Microsoft Excel” and “Excel” in the Google Play Store both display the target application. LlamaTouch extracts the content of the annotated textbox, and then compares the text with nodes in the matched screen using the same approach as in fuzzy screen info match.
3. Action match. Although the initiative of LlamaTouch is to detect state transfer during task execution. There are still cases for evaluating concrete actions performed on the screen, such as clicking one specific UI component, or typing the correct captcha. LlamaTouch directly compares actions and their parameters of the annotated ones with their counterparts on the matched screens.
4. System state match is usually more efficient and accurate than merely comparing UI states for specific tasks that involve deterministic system states, e.g., installed applications. LlamaTouch currently supports checking whether an application is installed or not. Such system state is recorded during task execution in real-world mobile devices ($\S$§3.1). LlamaTouch will check whether the annotated system state is identical to that of the last screen in the task execution trace.
Through the multi-level state matching algorithm, LlamaTouch achieves high evaluation accuracy on real-world task execution traces, while preserving the scalability of evaluating on static datasets. Experiments in $\S$§4 show the faithful evaluation of LlamaTouch.
4. Evaluation
4.1. Dataset
Table 3. LlamaTouch’s dataset statistics.
Task Category # Tasks # Apps Avg. Task Steps
Tasks from AITW (Rawles et al., 2023) 102 26 7.35 (2-19)
LlamaTouch Generated 393 46 5.67 (3-42)
Total 495 57 7.01 (2-42)
LlamaTouch uses a combination of tasks in the previous AITW dataset (Rawles et al., 2023) and self-constructed ones for evaluation. Table 3 shows the statistics of this dataset. We sampled 102 tasks from AITW (Rawles et al., 2023), one of the most popular datasets for mobile device control. By observing the limited number of evaluated applications, we generated 395 tasks in 46 applications. Overall, this dataset includes 495 tasks, covering 57 popular Android applications with different difficulty levels (measured by the number of steps to finish a task). Every single data entry consists of three types of data.
  • A task description in natural language, e.g., empty the shopping cart on bestbuy.
  • A sequence of raw UI representations when recording these tasks. Each raw UI representation contains a screenshot, VH, activity name, and action performed on the current screen.
  • The annotated essential states by human annotators.
4.2. Experiment Setup
Mobile environment. LlamaTouch primarily utilizes an x86-64 Android emulator (Google, 2023) as the mobile environment for task execution. The Android emulator is configured with Android 12 (API level 31) with Google Play Services deploying on a Linux server with Ubuntu 18.04 OS. Tasks involving applications that are not compatible with the Android emulator (e.g., Snapchat) will be tested on a real Google Pixel 5 device with the Android 14 OS.
Mobile agents. We selected four mobile agents that cover diverse types of brains, including supervised learning models, LLM, and large multi-modality models.
  • AutoUI (Zhan and Zhang, 2023) that employs a multimodal transformer model based on BLIP-2 (Li et al., 2023) and t5 (Raffel et al., 2020) with its variants as the brain for decision making. It takes pixel-level screenshots and task descriptions as the input.
  • AutoDroid (Wen et al., 2023) that uses LLMs for device control. It first takes textual screen VH and simplifies it to explicit, readable HTML representation. Then, the simplified HTML representation will be processed by LLMs and generate the corresponding action. We selected GPT-4 (gpt-4-0125-preview) (Achiam et al., 2023) as the backend of AutoDroid.
  • AppAgent (Yang et al., 2023) that utilizes GPT-4V (OpenAI, 2023) to comprehend the screenshots of mobile devices and then dispatch controlling commands to LlamaTouch. In our evaluation, we used the version AppAgent without document pre-exploration for simplicity.
  • CoCo-Agent (Ma et al., 2024) that uses LLaVA (Liu et al., 2023) as the brain. It is trained on a subset of the original AITW dataset.
All actions generated by these agents are applied to LlamaTouch’s Android environment through one adaption layer.
Evaluation methodologies. We used the following methodologies for evaluating whether an agent completes a task.
  • Step-wise action match is widely-utilized in evaluating mobile agents on well-established datasets (Rawles et al., 2023; Sun et al., 2022; Wen et al., 2023; Zhan and Zhang, 2023). It compares the agent-generated action sequences in real-world environments with ground-truth action sequences in the dataset. When two action sequences are identical, the task is treated as completed. We utilize the action match algorithm in AITW (Rawles et al., 2023): two actions are matched only if they have identical action types and parameters.
  • Longest common subsequence (LCS)-based action match is proposed in AndroidArena (Xing et al., 2024). It is built upon step-wise action match by tolerating redundant actions between ground-truth ones. Two action sequences are matched when the execution action sequence contains the ground-truth sequence as the subsequence.
  • LlamaTouch that compares agent execution traces with our annotated essential states to check whether a task is completed.
  • Human validates whether a task is completed based on the agent execution traces. Humans are instructed not only to focus on the action (and its parameter) on each UI representation, but also the whole application state transfer during task execution. The results of human validation are treated as ground truth for comparing the accuracy of the three evaluation approaches.
Metrics. We primarily compare the (1) end-to-end task completion rate (TCR) of different agents and (2) the accuracy of different evaluation approaches, taking human validation as the ground-truth result. In step-wise action match and LCS-based action match, a task is completed only when the two action sequences are identical or the ground-truth one serves as the subsequence of the agent-generated action sequence, respectively. In LlamaTouch, a task is treated as completed when the agent execution trace passes through all essential states. For accuracy, a task is accurately evaluated only if the evaluation result matches the human validation result. The accuracy of an evaluation approach denotes the proportion of tasks correctly evaluated out of all tasks.
4.3. Task Completion Rate and Accuracy
Table 4. End-to-end task completion rate (TCR %) and accuracy (Acc. %) of different evaluation approaches of all tasks.
Mobile Agent
Step-wise
action match
LCS action
match
LlamaTouch Human
TCR Acc. TCR Acc. TCR Acc. TCR
AutoUI 0.00 98.18 0.00 98.18 4.44 96.57 1.82
AutoDroid 0.00 85.98 0.00 85.98 14.84 91.87 14.02
AppAgent 0.00 93.33 0.61 93.13 10.91 94.95 6.67
CoCo-Agent 0.00 97.97 0.00 97.97 4.47 96.34 2.03
Average 0.00 93.86 0.15 93.81 8.67 94.93 6.14
Table 4 presents the end-to-end TCR and accuracy of different evaluation designs. All three approaches achieve more than 90% accuracy on average. However, the step-wise action match and LCS-based action match fall short in recognizing tasks correctly executed by mobile agents, resulting in nearly 0% TCR. The high accuracy rates of these two approaches are attributed to the large portion of incomplete tasks; only 6% of tasks are deemed completed upon human validation. Compared to action match on static datasets, the TCR evaluated by LlamaTouch is 8.67%, which closely aligns with the results of human validation.
To demonstrate LlamaTouch’s effectiveness in addressing the issue of false negative results that pervasively occurred in previous evaluation methods, we focused on tasks considered successfully completed by human evaluation, excluding all incomplete tasks. Table 5 displays the number of tasks completed by agents and the evaluation accuracy for these tasks. Overall, agents successfully completed 30 tasks on average. Among these tasks, both step-wise action match and LCS-based action match achieve no more than 0.1% evaluation accuracy. This indicates they are unable to faithfully evaluate tasks executed in real-world environments using static datasets. Notably, LlamaTouch exhibits an average accuracy of 79% in validating these task execution traces, significantly reducing the percentage of false negative cases observed with other evaluation methods. The findings uncover LlamaTouch’s proficiency in evaluating UI automation tasks under real-world settings.
Table 5. Accuracy (Acc. %) of different evaluation approaches among all successful tasks in human validation.
Mobile Agent
Step-wise
action match
LCS action
match
LlamaTouch Human
Acc. Acc. Acc. # success tasks
AutoUI 0.00 0.00 77.78 9
AutoDroid 0.00 0.00 73.91 69
AppAgent 0.00 3.03 93.94 33
CoCo-Agent 0.00 0.00 70.00 10
Average 0.00 0.76 78.91 30
$4.$ Ablation Study
Table 6. Ablation study on different essential state primitives.
Evaluation design All tasks AITW tasks LlamaTouch tasks
TCR Acc. TCR Acc. TCR Acc.
Complete LlamaTouch 8.67 94.93 17.46 89.43 6.38 96.36
LlamaTouch W/O exact match 86.77 18.85 82.55 29.75 87.87 16.02
+ activity exact match 23.21 81.60 41.05 68.29 18.58 85.06
+ action exact match 86.62 19.01 81.81 30.49 87.87 16.02
+ UI component exact match 15.86 88.15 40.30 68.55 9.51 93.24
+ system state exact match 85.41 20.22 75.91 36.40 87.87 16.02
LlamaTouch W/O fuzzy match 10.54 93.26 23.36 84.02 7.21 95.66
+ screen-level fuzzy match 10.24 93.36 22.13 84.76 7.15 95.60
+ textbox fuzzy match 8.97 94.83 18.69 88.69 6.45 96.43
LlamaTouch achieves high accuracy when evaluating agent execution traces in real-world environments, attributing to the annotation and implementation of different essential state primitives. In this section, we evaluate the effectiveness of two types of match design in LlamaTouch: fuzzy match and exact match. We aggregate the execution traces of all mobile agents evaluated in $\S$§4.3. As for evaluation implementation, we disable exact match and fuzzy match separately, and then enable different primitives once per time. We present the results on (1) all tasks; (2) tasks in AITW (Rawles et al., 2023); and (3) tasks generated in LlamaTouch. The results of task completion rate and accuracy are shown in Table 6.
Exact match ensures key information in two screens are identical. As shown in the results, exact match contributes a significant portion of evaluation accuracy improvement. For example, without all exact match primitives, the evaluation accuracy of LlamaTouch on all tasks drops from 95% to 19%. Among all exact match primitives, activity match highly improves the validation accuracy. This is not only attributed to its simple yet efficient ability to locate functional application screens, but also the large portion of annotated activity primitives, i.e., 35% out of all primitives in our dataset. UI component match is also necessary for faithful evaluation in LlamaTouch, as it is typically with screen location primitives such as activity to detect critical information within the matched screen. Action and system state exact match only improve a little accuracy as only a few tasks are annotated with these primitives.
Fuzzy match does not show notable improvement in evaluation accuracy, although there are 30 screen-level fuzzy match primitives and 52 textbox fuzzy match primitives in the dataset. Seldom of tasks with these primitives can be completed by the four agents we evaluated, leading to a slight improvement. Although primitives about fuzzy match are not well explored, they play an important role in dealing with screen content or UI layout dynamics in the real world. With more performant mobile agents in the future, they will be further explored and evaluated.
4.5. Absolute Capabilities of Mobile Agents
In this section, we present the absolute capabilities of different mobile agents in mobile UI automation tasks. We first categorize tasks according to their sources: AITW and our self-constructed dataset. This aims to reveal whether an agent that has previously learned on a well-established dataset has the capability to adopt new applications/tasks. Second, we categorize tasks according to their difficulty, measured by the number of steps required to complete a task in the ground-truth dataset. Tasks are further divided into three difficulty levels: easy (steps $\leq$ 4), medium (4 <<< steps $\leq$ 8), and high (step >>> 8).
Table 7. The comparison of task completion rate categorized by task sources (e.g., from the previous dataset AITW or from LlamaTouch’s generated dataset).
Agent End-to-end TCR
Overall AITW Generated
AutoUI 4.44 12.75 2.29
AutoDroid 14.84 22.77 12.79
AppAgent 10.91 21.57 8.14
CoCo-Agent 4.47 12.75 2.31
Performance among datasets. Among the four evaluated agents, AutoUI and CoCo-Agent are previously trained on AITW. AutoDroid and AppAgent directly invoke GPT-4 and GPT-4V, respectively. We separately show end-to-end TCR of all agents on tasks in AITW and LlamaTouch’s generated tasks to show their performance generalization to new scenarios (i.e., new tasks and new applications). Results in Table 7 show that AutoDroid and AppAgent also show higher TCR in AITW tasks than LlamaTouch’s generated tasks. This mainly attributes to the task complexity gap between the two datasets. AutoUI and CoCo-Agent achieve 12.75% TCR on tasks in AITW. However, for generated tasks in LlamaTouch, both of them operate poorly: with only 2.29% and 2.31% TCR, respectively. The TCR gap between datasets reveals their lack of capability to adapt to unseen tasks before. Considering massive real-world applications, mobile agents with strong generalization capabilities to unseen scenarios are more competitive.
Performance under different task complexity. We categorize tasks into three difficulty levels according to the step number required to complete the task in the datasets. The results are shown in Table 8. Generally, mobile agents can better complete simple tasks that require fewer steps. There is a significant drop in TCR when task complexity increases. AutoDroid outperforms the other three agents for tasks under all difficulties. AppAgent achieves a similar TCR among all tasks as AutoDroid. We deem that the knowledgable GPT-4/4V can better interpret task descriptions in natural language and pixel-level or textual UI representations. Therefore, agents based on GPT-4/4V achieve high TCR. However, there is still huge space for mobile agents to improve their capabilities in UI automation tasks.
Table 8. The comparison of task completion rate between mobile agents. Tasks are categorized to different difficulty levels according to the number of steps required to finished it in the ground-truth dataset.
Agent End-to-end TCR
[…]
AutoUI 4.44 4.95 4.19 4.82
AutoDroid 14.84 27.00 12.94 7.32
AppAgent 10.91 16.83 10.97 3.61
CoCo-Agent 4.47 7.92 3.91 2.41
5. Limitations of LlamaTouch
Supporting WebView-based apps. One major limitation of LlamaTouch is that its evaluation process requires Android’s view hierarchy to empower approximate UI layout matching and accurate UI component matching ($\S$§3.2). While some Android applications built with WebView (Google, 2024a) are unable to access their view hierarchy, making LlamaTouch cannot evaluate agent performance on these applications. However, only a minor number of applications are built with WebView, inspiring LlamaTouch’s solution in utilizing view hierarchy. When evaluating WebView-based apps are required, more advantaged techniques, such as screen similarity detection (Wu et al., 2023b, a), OCR/model-powered screen element recognition (Bunian et al., 2021; Xie et al., 2020), should be incorporated to retrofit LlamaTouch. We will consider this in future work.
Biases in identifying essential states. LlamaTouch requires human or other autonomous agents, e.g., GPT-4V, to identify and annotate essential states on pre-defined task execution traces. This process, however, could involve biases and leads to inaccuracy due to potentially limited knowledge in application execution and unseen circumstances. For example, given a task “Delete YouTube in the Google Play Store”, the pre-defined traces might explicitly navigate to the app page of “YouTube” and click the “Uninstall” button, which might be annotated as one essential state during task execution. However, they may ignore that this application might be not installed at all. In this case, there will be no explicit action of “clicking the uninstall button” while this task is still completed. Such biases might result in low accuracy in checking task completion rates. However, we think this limitation could be eliminated and refined by involving expert reviewing. Furthermore, for every single task description, involving diverse essential states on different potential task execution paths could better enhance the robustness of LlamaTouch’s evaluation design.
6. Conclusion
In this work, we proposed LlamaTouch, the first testbed for evaluating mobile agents with faithfulness and scalability in UI automation tasks. LlamaTouch enables mobile agents to be tested on real mobile devices. At the evaluation stage, it only matches the task execution traces with annotated essential states. LlamaTouch well tolerates different task execution paths and dynamic execution environments, significantly reducing false negative results that occurred in previous evaluation approaches. It achieves high evaluation accuracy like human validation, while preserving the scalability of evaluating on static datasets. By putting task execution on real mobile devices, we also reveal the limited capabilities of current mobile agents in UI automation tasks.
Appendix A Appendix
A.1. APIs Provided by LlamaTouch
LlamaTouch provides three categories of APIs for mobile agent integration and UI automation task execution. APIs, parameters, and their return values are listed in Table 9.
Table 9. APIs provided by LlamaTouch to mobile agents for UI automation task execution on real mobile devices.
API Category API Parameter Return Value
Metadata query get_task_instruction None str: A task description in natural language.
UI state query get_screenshot None
str: A base64 encoded string
representing the current screenshot.
get_view_hierarchy None
str: A string representing the
textual view hierarchy in XML.
Action space post_task_complete None None
post_task_impossible None None
post_press_home None None
post_press_back None None
post_click
x, y: Normalized x/y coordinates for a
targeted screen position.
None
post_type str: The text to be input. None
post_swipe
touch_x, touch_y: Normalized x/y
coordinates where the swipe begins.
lift_x, lift_y: Normalized x and y
coordinates where the swipe ends.
duration: The interval of the swipe gesture.
None